# Reduced Order Hysteretic Magnetization Model for Composite Superconductors

### v1

@@ Abstract
In this paper, we propose the Reduced Order Hysteretic Magnetization (ROHM) model to describe the magnetization and instantaneous power loss of composite superconductors. Once the parameters of the ROHM model are fixed based on reference simulations, it allows to directly compute the macroscopic response of composite superconductors without having to solve for the detailed current density distribution. It can be used as a homogenized model in large-scale superconducting systems in order to significantly reduce the computational work compared to detailed simulations. In this contribution, we focus on the case of a strand with twisted superconducting filaments subject to an external transverse magnetic field. We propose two levels of ROHM models: a rate-independent model that reproduces pure hysteresis without dynamic effects, and a rate-dependent model that generalizes the former by also reproducing dynamic effects observed in superconducting strands due to coupling and eddy currents. We describe the implementation and inclusion of these models in a finite element framework, discuss their parameter identification and finally demonstrate the capabilities of the approach in terms of accuracy and efficiency.
  • September 20, 2024
@@
*                                    This work is a preliminary version. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@
*                                    This work is a preliminary version. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Large-scale superconducting systems are multi-scale structures made of a high number of wounded superconducting cables, which themselves consist of multifilamentary strands or complex arrangements of tapes. The magnetic response of these systems is hysteretic and rate-dependent. Hysteresis is created by the irreversible motion of flux vortices among pinning centers in superconducting parts [1, 2, 3, 4], whereas rate-dependency is caused by eddy currents, and coupling current dynamics between different conductors [5, 6, 7].
Hysteresis, eddy current, and coupling current effects induce loss in transient regimes [5] as well as field distortions [8]. Their computation is therefore crucial, e.g., for computing the load on the cryogenic system, the temperature and stability margins, field errors, or for the design of quench protection devices.
@@
Besides these advantages, the power law introduces a strong nonlinearity in the equations, which makes the numerical models computationally demanding to solve, already for single strands or cables [11]. As a consequence, modelling large-scale superconducting systems in all their details down to the conductor level is completely unrealistic in practice. Faster, more efficient, methods are necessary.
@@
Such techniques have been extensively studied in non-superconducting electromagnetic systems, such as periodical structures [13, 14, 15] with bundles of wires of any shape, described in the frequency or time domain [16, 17]. They rely on the design of equivalent conductivity and permeability parameters, that must be fitted on reference solutions, or on experimental measurement results.
Following a similar approach for superconducting systems requires to use dedicated history-dependent parameters in order to reproduce their hysteretic response.
A few hysteresis models have been proposed for modelling superconductor. A vector model is proposed in [8, 18] for field quality computation in accelerator magnets. This model describes the magnetization of superconducting filaments using nested magnetization ellipses. Another hysteresis model based on the Preisach model [19] was proposed in [20, 21] in the form of an equivalent circuit element. Yet another vector model based on a variational approach was described in [22]. These models however only describe rate-independent hysteresis and are therefore not suited for modelling transient effects in composite superconductors.
@@
The energy-based model offers several advantages. First, it gives a direct access to instantaneous dissipated and stored power by conveniently separating the magnetic field into several contributions. Second, it offers a high flexibility and modularity thanks to an approach involving several sub-elements that can be combined. Third, it is consistently defined as a vector, and not scalar, model from the beginning [27]. Fourth, it can be directly generalized to include rate-dependent effects [28]. Finally, the equations of this model consist in simple explicit tests, which lead to straightforward implementations and very efficient simulations.
In this paper, we adapt and extend the state-of-the-art energy-based model to composite superconductor hysteresis by focussing on the case of a superconducting multifilamentary strand subject to a transverse magnetic field, i.e., a field perpendicular to the strand axis. This situation is relevant for most applications and constitutes a first step towards the homogenization of a complete magnet winding. This is also a challenging problem since, in addition to the filament hysteresis effect, it also involves two types of rate-dependent effects: eddy currents in normal conductors and interfilamentary coupling currents. For this paper, we assume that the strand carries no transport current. The inclusion of transport current will be the focus of further work. We also consider a constant temperature and do not model thermal effects.
We propose two models: (i) a rate-independent model which describes the superconducting hysteresis, and (ii) a rate-dependent model, which generalizes the first model by also describing the eddy current and coupling current effects, in addition to the superconducting hysteresis.
The models are general and can be adapted to a variety of composite superconductors, such as strands, tapes, or cables. They can also be easily implemented in any FE framework allowing for user-defined material properties.
This paper is structured as follows. In Section 2, we introduce the model of a multifilamentary strand and describe its response in terms of magnetization and loss for transverse fields in a range of frequencies and amplitudes. In Section 3, we present the equations leading to rate-independent and rate-dependent hysteresis models. In Section 4, we discuss the implementation of the equations and describe how to include them in a two-dimensional (2D) FE framework with a $\phi$-formulation. We then propose in Section 5 a parameter identification procedure for the models. Finally, in Sections 6 and 7, we apply the hysteresis models in rate-independent and rate-dependent situations, respectively, and demonstrate their potential in the context of homogenization techniques.
@@ 2 Multifilamentary Strand Dynamics
In this section, we present the response of a composite multifilamentary superconducting strand subject to a transverse magnetic field in terms of power loss and magnetization. To this end, we model the detailed current density distribution inside the strand with a FE method and we analyze the different dynamics that result from the strand composite structure.
@@ 2.1 Problem definition
We consider a strand of diameter $d$ represented in Fig. 1. It consists of $N_{\text{f}}$ Nb-Ti filaments twisted with twist pitch length $p$ and embedded in a conducting copper (Cu) matrix. The geometrical parameters of the strand are summarized in Table 1. The strand carries no transport current but is subject to an applied (app) transverse magnetic field $\boldsymbol{h}_{\text{app}}(t)=h_{\text{max}}\sin(2\pi ft)\hat{\boldsymbol{e}}_{y}$ of amplitude $h_{\text{max}}$ (A/m), frequency $f$ (Hz), and direction $\hat{\boldsymbol{e}}_{y}$.
Figure 1: Composite strand with Nb-Ti filaments embedded and twisted in a copper matrix (cylindrical outline) subject to a time-varying transverse magnetic field $\boldsymbol{h}_{\text{app}}(t)$.
Number of filaments ($N_{\text{f}}$) 54
Filament diameter $90~{}\upmu$m
Filament center-to-center spacing $110~{}\upmu$m
Strand diameter ($d$) $1$ mm
Twist pitch length ($p$) $19$ mm
Table 1: Geometrical parameters of the strand.
We fix the temperature to $1.9$ K. The resistivity of the Nb-Ti filaments is described by the power law [9] with power index $n=30$ and the field-dependent critical current density $j_{\text{c}}(\boldsymbol{b})$ (A/m2) given in Fig. 2, with $\boldsymbol{b}$ (T) the local magnetic flux density, defined as $\boldsymbol{b}=\mu_{0}\boldsymbol{h}$, with $\mu_{0}=4\pi\times 10^{-7}$ H/m the permeability of vacuum. The resistivity of the copper matrix, accounting for magneto-resistance, is obtained from the STEAM material library [30], with a residual resistivity ratio $\text{RRR}=100$.
Figure 2: Field-dependent critical current density $j_{\text{c}}(\boldsymbol{b})$. The maximum value is $j_{\text{c}}(\boldsymbol{0})=2.4\times 10^{10}$ A/m2.
We model the strand response with the CATI method, proposed in [31]. This method accounts for 3D effects due to the strand twist by using a pair of 2D models solved on a cross section of the strand, and coupled via circuit equations. As it is based on 2D models, it offers fast simulations. The method is implemented in GetDP [32] within FiQuS [33]. FiQuS is developed at CERN as part of the STEAM framework [34]. Geometry and mesh are performed by Gmsh [35]. All the software is open-source and free to use. The code is available online [^2] https://gitlab.cern.ch/steam/analyses/cati-strand.
2.2 Magnetization and loss
We focus on magnetization and power loss. These are the two quantities we want to reproduce directly with the hysteresis model without having to compute the detailed current density distribution.
@@ 2.2 Magnetization and loss
with $a=\pi d^{2}/4$ the surface area of the strand cross section $S$, $\boldsymbol{x}$ the position vector, and $\boldsymbol{j}$ the current density. The multiplication by $2$ accounts for current loops closing at infinity in this 2D problem, as justified in A.
The instantaneous power loss per unit length $P_{\text{tot}}$ (in W/m) is obtained by integrating the instantaneous power loss density, or Joule loss, over the strand cross section, including the Nb-Ti filaments and the Cu matrix,
@@
with $\boldsymbol{e}=\rho\boldsymbol{j}$ the electric field (and $\rho$ the resistivity). The power loss per cycle and per unit length (in J/m) is evaluated as
[…]
To make the interpretation of the loss easier, and also because it will be helpful for the construction of the hysteresis model, the total power loss per cycle $Q_{\text{tot}}$ is decomposed in three contributions, as was done in [31]: hysteresis, coupling, and eddy loss. The hysteresis loss is the loss induced by currents flowing in the superconducting filaments. The coupling loss is the loss induced by currents flowing in the conducting matrix, perpendicular to $\hat{\boldsymbol{e}}_{z}$ (the coupling currents). Finally, the eddy loss is the loss induced by currents flowing in the conducting matrix along $\hat{\boldsymbol{e}}_{z}$.
Magnetization and loss are closely related to each other. Indeed, the area inside a closed magnetization loop is proportional to the total loss per cycle [37, 38]. We have, in terms of the applied magnetic field $\boldsymbol{h}_{\text{app}}$,
@@ 2.3 Dynamic response of the strand
The strand response is computed by the CATI method for a range of field amplitudes $h_{\text{max}}$ and frequencies $f$. The power loss per cycle induced by the transverse field is represented in Fig. 3 as a function of $f$ for three different values of $h_{\text{max}}$. The $y$-component of the magnetization vector, $m$, is represented in Figs. 4 and 5 for $\mu_{0}h_{\text{max}}=2$ T as a function of $\mu_{0}h_{\text{app}}(t)$ for selected frequencies.
@@
At low frequencies, when the rate of change of the transverse magnetic field is not high enough for significant coupling currents and eddy currents to take place in the matrix, the superconducting filaments behave close to uncoupled. In that regime, most loss is due to hysteresis in the filaments. A typical magnetization loop in that case is represented in Fig. 4. The magnetization decreases when the field increases as a result of the field-dependent critical current density. In this regime, a good first approximation is to consider that the response is rate-independent (it is not strictly the case in reality due to the finite value of the $n$ index in the power law). To this end, a rate-independent hysteresis model can be used.
Figure 4: Strand magnetization along $\hat{\boldsymbol{e}}_{y}$ for $\mu_{0}h_{\text{max}}=2$ T and $f=0.01$ Hz. Same result as the first curve in Fig. 5.
When the frequency increases, the response of the strand becomes strongly rate-dependent. As shown in Fig. 3, coupling and eddy loss can dominate the total loss, and the filament loss is not constant with frequency. As a result, the loss per cycle can exhibit more than one order of magnitude of variation as a function of frequency for an identical applied field amplitude. Similarly, the magnetization strongly depends on the frequency, as shown in Fig. 5. To reproduce these regimes, there is a clear need for a rate-dependent hysteresis model. The design of such a model benefits from a good understanding of the physics of the curves in Figs. 3 and 5.
Figure 5: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T and for different frequencies $f$. Results from [31] using Eq. (1).
For low field amplitudes, the dynamics of the coupling currents and of the associated loss is well described by analytical models [5, 39]. The coupling loss follows follows a typical bell curve with a maximum at a characteristic frequency $f_{\text{c}}$.
With increasing field amplitude, the peak of the coupling loss curve shifts towards lower frequencies, as can be seen in Fig. 3 for $\mu_{0}h_{\text{max}}=2$ T. This is associated with the saturation of superconducting filaments, which limit the coupling currents amplitude, and to a change of the effective permeability, due to the loss of diamagnetic effect due to saturated filaments, as discussed in [5].
The coupling currents also give rise to a change of regime for the filament magnetization and loss. As mentioned previously, at low frequencies, $f\lesssim f_{\text{c}}$, they are mostly uncoupled, producing a relatively small magnetization. On the other hand, for $f\gtrsim f_{\text{c}}$, they are coupled and exchange currents via coupling currents. As a result, their magnetization is increased, see for example the situation for $f=10$ Hz in Fig. 5. Whether the filament loss for coupled filaments is higher or lower than that for uncoupled filaments depends on $\mu_{0}h_{\text{max}}$, as can be seen in Fig. 3.
At even higher frequencies, eddy currents in the conducting matrix become the dominating factor for loss and magnetization. As a result of eddy currents and the skin effect, the inner part of the strand, containing the filaments, is shielded from the outer field and the filament loss decreases accordingly. In that regime, the magnetization curve approaches the shape of an ellipse [37] that gets thinner and thinner with increasing frequency, and of slope approaching $-2$ because of demagnetization effects (this is further explained in Section 3.1[38].
@@ 3 Reduced Order Hysteretic Magnetization Model
The simulation time for each of the simulations presented in the previous section is of the order of one hour with the CATI method. This is fast and convenient enough to get a good understanding of a single strand response. However, for simulating a full-scale magnet, containing thousands of strands, one cannot afford such a detailed description of the current density at the strand level. Instead, one has to use an alternative model that produces similar macroscopic response in terms of power loss and magnetization, which are both history-dependent and rate-dependent, but without requiring a detailed calculation of the current density, so as to strongly reduce the computational cost. This idea is illustrated in Fig. 6. We present such a model in this section, the Reduced Order Hysteretic Magnetization (ROHM) model.

### v2

@@ Abstract
In this paper, we propose the Reduced Order Hysteretic Magnetization (ROHM) model to describe the magnetization and instantaneous power loss of composite superconductors subject to time-varying magnetic fields. Once the parameters of the ROHM model are fixed based on reference simulations, it allows to directly compute the macroscopic response of composite superconductors without the need to solve the detailed current density distribution. It can then be used as part of a homogenization method in large-scale superconducting models to significantly reduce the computational effort compared to detailed simulations. In this contribution, we focus on the case of a strand with twisted superconducting filaments subject to a time-varying transverse magnetic field. We propose two levels of ROHM models: (i) a rate-independent model that reproduces hysteresis in the filaments, and (ii) a rate-dependent model that generalizes the first level by also reproducing dynamic effects due to coupling and eddy currents. We describe the implementation and inclusion of these models in a finite element framework, discuss how to deduce the model parameters, and finally demonstrate the capabilities of the approach in terms of accuracy and efficiency over a wide range of excitation frequencies and amplitudes.
  • October 25, 2024
@@
*                                    This work has been submitted to a journal for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@
*                                    This work has been submitted to a journal for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Large-scale superconducting magnets are multi-scale structures made of a high number of wound superconducting cables, which themselves consist of multifilamentary strands or complex arrangements of tapes. The magnetic response of these systems is hysteretic and rate-dependent. Hysteresis is created by the irreversible motion of flux vortices among pinning centers in superconducting elements [1, 2, 3, 4], whereas rate-dependency is caused by eddy and coupling current dynamics [5, 6, 7].
Hysteresis, eddy current, and coupling current effects induce loss in transient regimes [5] as well as field distortions [8]. Their accurate description and computation is therefore crucial, e.g., for computing the load on the cryogenic system, the temperature and stability margins, field errors, or for the design of quench protection devices.
@@
Besides these advantages, the power law introduces a strong nonlinearity in the equations, which makes the numerical models computationally demanding to solve, already for single strands or cables [11]. As a consequence, modelling large-scale superconducting systems in all their details down to the superconductor level is completely unrealistic in practice. A faster and more efficient method is necessary.
@@
Such techniques have been extensively studied in non-superconducting electromagnetic systems, such as periodical structures with bundles of wires of any shape [13, 14, 15], described in the frequency or time domain [16, 17]. They rely on the design of equivalent conductivity and permeability parameters, that must be fitted on reference solutions.
Following a similar approach for superconducting systems requires the use of dedicated history-dependent parameters in order to reproduce their hysteretic response.
A few hysteresis models have been proposed for modelling superconductors. A vector model is proposed in [8, 18] for field quality computation in accelerator magnets. This model describes the magnetization of superconducting filaments using nested magnetization ellipses. Another hysteresis model based on the Preisach model [19] was proposed in [20, 21] in the form of an equivalent circuit element. Yet another vector model based on a variational approach was described in [22]. These models however only describe rate-independent hysteresis and are therefore not suited for modelling transient effects in composite superconductors.
@@
The energy-based model offers several advantages. First, it gives a direct access to instantaneous dissipated and stored power by conveniently separating the magnetic field into several contributions. Second, it offers a high flexibility and modularity thanks to an approach involving several sub-elements that are combined to form the final model. Third, it is consistently defined as a vector, and not scalar model from the beginning [27], and hence applies to two-dimensional (2D) or three-dimensional (3D) settings. Fourth, it can be directly generalized to include rate-dependent effects [28]. Finally, the equations of this model consist in simple explicit tests, which leads to straightforward implementations and very efficient simulations.
In this paper, we adapt and extend the state-of-the-art energy-based model to composite superconductor hysteresis by focussing on the case of a superconducting multifilamentary strand subject to a transverse magnetic field, i.e., a field perpendicular to the strand axis. This situation is relevant for most applications and constitutes a first step towards the homogenization of a complete magnet winding. This is also a challenging problem since, in addition to the filament hysteresis effect, it also involves two types of rate-dependent effects, eddy currents and interfilamentary coupling currents in the strand matrix, as well as filament hysteresis change due to coupling and eddy currents. For this paper, we assume that the strand carries no transport current. The inclusion of transport current will be the focus of further work. We also consider a constant temperature and do not model thermal effects.
We propose the Reduced Order Hysteretic Magnetization (ROHM) model, and present two variations of it: (i) a rate-independent model which describes the superconducting hysteresis, and (ii) a rate-dependent model, which generalizes the former by also describing eddy current and coupling current effects, in addition to the superconducting hysteresis.
The ROHM model is general and can be adapted to a variety of superconductors or composite superconductors, such as strands, tapes, stacks of tapes, cables, or bulks. It can also be easily implemented in any FE framework that allows for user-defined material properties.
This paper is structured as follows. In Section 2, we describe the response of a multifilamentary strand in terms of magnetization and loss for transverse fields with a range of frequencies and amplitudes, in order to motivate the design of the ROHM model. In Section 3, we present the equations defining the rate-independent and rate-dependent ROHM models. In Section 4, we discuss the implementation of these equations and we describe how to include them in a 2D FE framework with a $\phi$-formulation. We then propose in Section 5 a parameter identification approach for fixing the ROHM model parameters based on reference solutions. Finally, in Sections 6 and 7, we apply the ROHM model in rate-independent and rate-dependent situations, respectively, and demonstrate its potential in the context of homogenization techniques.
@@ 2 Multifilamentary Strand Dynamics
In this section, we describe the response of a composite superconducting strand subject to a transverse magnetic field in terms of power loss and magnetization. To this end, we model the detailed current density distribution inside the strand with a FE method, the CATI method [30] implemented in FiQuS [31, 32, 33], and we analyze the different dynamics resulting from the strand composite structure.
@@ 2.1 Problem definition
We consider a multifilamentary strand such as in Fig. 1. It consists of twisted superconducting filaments embedded in a cylindrical conducting matrix. The strand carries no transport current but is subject to a time-varying transverse magnetic field $\boldsymbol{h}_{\text{app}}(t)=h_{\text{max}}\sin(2\pi ft)\hat{\boldsymbol{e}}_{y}$ of amplitude $h_{\text{max}}$ (A/m) and frequency $f$ (Hz).
Figure 1: Composite strand with twisted superconducting filaments embedded in a conducting matrix (cylindrical outline) subject to a time-varying transverse magnetic field $\boldsymbol{h}_{\text{app}}(t)$.
2.2 Magnetization and power loss
We focus on magnetization and power loss. These are the two quantities to be reproduced with the ROHM model.
@@ 2.2 Magnetization and power loss
with $a$ the surface area of the strand cross section $S$, $\boldsymbol{x}$ the position vector, and $\boldsymbol{j}$ the current density. The multiplication by $2$ accounts for current loops closing at infinity, as justified in A.
The instantaneous power loss per unit length $P_{\text{tot}}$ (W/m) is obtained by integrating the instantaneous power loss density, or Joule loss density, over the strand cross section, including the filaments and the matrix,
@@
with $\boldsymbol{e}=\rho\boldsymbol{j}$ the electric field and $\rho$ the resistivity. The energy loss per cycle and per unit length (J/m) is evaluated as
$Q_{\text{tot}}=\int_{1/f}^{2/f}P_{\text{tot}}(t)\,\text{d}t.$ (3)
To make the interpretation of the loss easier, and to aid the construction of the ROHM model, the total energy loss per cycle $Q_{\text{tot}}$ is decomposed in three contributions: hysteresis, coupling, and eddy loss. The hysteresis loss is the loss induced by currents flowing in the superconducting filaments. The coupling loss is the loss induced by currents flowing in the conducting matrix perpendicular to $\hat{\boldsymbol{e}}_{z}$ (the coupling currents). Finally, the eddy loss is the loss induced by currents flowing in the conducting matrix along $\hat{\boldsymbol{e}}_{z}$.
Magnetization and loss are related to each other. Indeed, the area inside a closed magnetization loop is proportional to the total loss per cycle [35, 36]. We have, in terms of the applied magnetic field $\boldsymbol{h}_{\text{app}}$,
@@ 2.3 Dynamic response of the strand
We now describe the strand response as a function of the field amplitude and frequency, and highlight its hysteretic behavior as well as the associated rate-dependency.
For illustration, we consider a strand made of $54$ Nb-Ti filaments (diameter $90$ $\upmu$m) twisted with a twist pitch length of $19$ mm and embedded in a Cu matrix with a diameter of $1$ mm (same geometry as in [30]). We fix the temperature to $1.9$ K. The resistivity of the Nb-Ti filaments is described by the power law [9] with power index $n=30$ and a field-dependent critical current density $j_{\text{c}}(\boldsymbol{b})$ [37], with $\boldsymbol{b}$ (T) the local magnetic flux density, defined as $\boldsymbol{b}=\mu_{0}\boldsymbol{h}$, with $\mu_{0}=4\pi\times 10^{-7}$ H/m. The resistivity of the copper matrix accounts for magneto-resistance, with a residual resistivity ratio $\text{RRR}=100$ at zero field [37].
2.3.1 Rate-independent regime
At low frequencies (here for $f\lesssim 0.01$ Hz), the rate of change of the transverse magnetic field generates negligible coupling and eddy currents in the matrix. In this regime, the superconducting filaments behave almost uncoupled and most of the loss is due to hysteresis in them. A typical magnetization loop is shown in Fig. 2. The magnetization decreases when the field increases as a result of the field-dependent critical current density.
Strictly speaking, the finite value of $n$ creates a small rate-dependency, but in good approximation, the response can be considered rate-independent. A rate-independent ROHM model can therefore be used to describe the strand response in this regime.
Figure 2: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T, $f=0.01$ Hz.
2.3.2 Rate-dependent regime
When the frequency increases, the response of the strand is no longer rate-independent. As shown in Fig. 3, coupling and eddy loss can no longer be neglected, and the hysteresis loss varies significantly with frequency. As a result, the total loss per cycle is strongly rate-dependent. The magnetization also strongly depends on the frequency, as shown in Fig. 4, where the $y$-component of the magnetization vector, $m$, is shown as a function of $\mu_{0}h_{\text{app}}(t)$ for different frequencies.
To reproduce magnetization and loss in this regime, there is a clear need for a rate-dependent ROHM model. The design of such a model benefits from a good understanding of the physics behind the curves in Figs. 3 and 4, which we briefly summarize below.
@@ 2.3.2 Rate-dependent regime
Figure 4: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T and for different frequencies $f$. Results from [30] using Eq. (1). The curve for $f=0.01$ Hz is shown in more detail in Fig. 2.
For low field amplitudes, the dynamics of the coupling currents and of the associated loss is well described by analytical models [5, 38]. The coupling loss follows a typical bell curve with a maximum at a frequency $f_{\text{c}}$.
With increasing field amplitude, the peak of the coupling loss curve shifts towards lower frequencies, as can be seen in Fig. 3 for $\mu_{0}h_{\text{max}}=2$ T. This is associated with the saturation of superconducting filaments, which limits the coupling currents amplitude, and to the loss of the diamagnetic effect in saturated filaments, as discussed in [5].
The coupling current dynamics influences the filament magnetization and hysteresis loss. At low frequencies, $f\lesssim f_{\text{c}}$, filaments are mostly uncoupled, whereas at higher frequencies, $f\gtrsim f_{\text{c}}$, they exchange currents via coupling currents. As a result, the overall magnetization increases, see for example the situation for $f=10$ Hz in Fig. 4, and the hysteresis loss is affected. Whether the hysteresis loss per cycle for coupled filaments is higher or lower than that for uncoupled filaments depends on the applied field amplitude, as can be seen in Fig. 3.
At even higher frequencies, eddy currents in the conducting matrix become the dominating factor for loss and magnetization. As a result of eddy currents and the skin effect, the inner part of the strand, containing the filaments, is shielded from the external field and the hysteresis loss decreases accordingly. In that regime, the magnetization curve approaches the shape of an ellipse [35] that becomes thinner and thinner with increasing frequency, and its slope approaches $-2$ because of demagnetization effects [36].
@@ 3 Reduced Order Hysteretic Magnetization Model
The simulation time for each of the simulations presented in the previous section is of the order of one hour with the CATI method [30]. This is fast and convenient enough to get a good understanding of a single strand response with a detailed description of the current density distribution (as discussed in [30] or shown later in Fig. 27). However, to simulate transients in a full-scale magnet containing thousands of strands, one cannot afford models involving such a detailed description at the strand level.
Instead, one can use an alternative model that produces accurate macroscopic response in terms of power loss and magnetization, which are both history-dependent and rate-dependent, but that does not rely on a detailed calculation of the current density. We present such a model in this section: the Reduced Order Hysteretic Magnetization (ROHM) model. This model is adapted from the energy-based model for ferromagnetic materials [27, 28].
We start in Section 3.1 by clarifying how the ROHM model can be used to model the strand. We then define the building blocks, referred to as cells, that will be combined for the construction of the ROHM model. Each cell defines a local hysteretic relationship between the magnetic field $\boldsymbol{h}$ and the magnetic flux density $\boldsymbol{b}$. We present in Section 3.2 a rate-independent cell, that we call the superconductor cell (S cell). In Section 3.3, we add contributions to account for eddy current and coupling current effects, resulting in a rate-dependent cell, that we call the composite superconductor cell (CS cell). Finally, we show in Section 3.4 how to combine several cells into a chain of cells, defining a complete ROHM model.
3.1 Homogenized model and internal field